Answer and do the following calculations:

1. Suppose your professor handed you a test tube with 2.0 mL of an E. coli broth culture in it and told you to make a 10–2 dilution of the entire culture. Explain how you would do this. Show your calculations.

2. How would you produce a 10–1 dilution of a 3 mL bacterial sample using the entire 3 mL volume?

3. You have 0.05 mL of an undiluted culture at a concentration of 3.6 X 106
CFU/mL. You then add 4.95 mL sterile diluent.
What is the dilution, and what is the final density of cells?

4. What is the purpose of serial dilutions?

5. Plating 1.0 mL of a sample diluted by a factor of 10–3 produced 43 colonies. What was the original cell
density in the sample?

Remember to use the following formulas for the above questions:

Calculate the original density in CFU/mL using the following formula:
OCD = CFU/ original sample volume

OCD = original cell density

CFU = colony forming units

and

V1D1 = V2D2

Note: unless told otherwise, the original dilution (D1) is always 1.
Answer #1

1. Dilution = initial volume / final volume

Initial volume = volume of culture = 2 mL

Final volume = initial volume + volume of diluent

10^-2 = 2 / final volume

Final volume = 2 / 10^-2 = 200 mL

Volume of diluent = 200 - 2 = 198 mL

To 2 mL culture, add 198 mL diluent.

2. 10^-1 = 3 / final volume

Final volume = 3 / 10^-1 = 30 mL

Volume of diluent = 30 - 3 = 27 mL

To 3 mL culture, add 27 mL diluent.

As per the guidelines, only 1st question can be answered.

Please rate high.